Convert the values of Kc to Kp Cl2(g)+Br2(g)=2BrCl(g) K =4.7x10-2 at 24.7 °C
Answer #1

T= 24.7 oC

= (24.7+273) K

= 297.7 K

Δ n = number of gaseous molecule in product - number of gaseous molecule in reactant

Δ n = 0

use:

Kp= Kc (RT)^Δ n

Kp = 4.7*10^-2*(0.08206*297.7)^(0)

Kp = 4.7*10^-2

Answer: 4.7*10^-2